A⃗ =( 26 m )x^+( -11 m )y^.

B⃗ =( 3.5 m )x^+( 16 m )y^.

Find the direction of vector A + B
Answer #1

m À = (26 å - 11 ŷ ) m B = (3.5 ñ +16 tibg à B = (26 +3.5) ê t (+1+16) 29.5 êts direction 6) tant s २०.5 9.6199 from tx tx an